OpenStudy (anonymous):

A balloon holds 26.6 kg of helium. What is the volume of the balloon if the pressure is 0.96 atm and the temperature is 23 °C?

OpenStudy (mathmate):

Hints: Ideal Gas Law: PV=nRT P=pressure in pascals (Pa). 101325 Pa = 1 atmosphere V=volume in m^3 (1 m^3=1000 L) T=temperature in degrees Kelvin, = degrees C + 273.15 n=number of moles, = m/M (m=mass of gas, M=molecular mass of helium=4.003 g) R=universal gas constant = 8.314 J/deg.K/mol Put it all together and solve for V (in m^3).